2. Valle de Estós

’POSETS’ PEAK 6 hours 15 mins. total time 4 hours (approx.) from the Forcau refuge 3,375 m. high

To reach the top of the Posets, it is recommended that you spend the night in one of the mountain refuges you will find on the way up, such as Estós or Angel Orús (being these the only ones opened in the Posets area in this time of the year). You also have to take with you the necessary material for snow and ice.

The arresting views from this peak are probably the best ones of the whole valley. Depart from the Angel Orús refuge (2,095 m.) in the Ball de Gris (see map). Go

up by a well-traced track to the Llardaneta ravine. Pass through a plleta to get into the spacious prairies area that will take you to the Gorgues de Llardaneta (2,475 m.). Turn left by going up the Canal Fonda (Deep Channel), usually covered in snow, up to the Collada del Diente (3,000 m.). Turn right again towards the crest which, although wide at the beginning, will soon narrow to become a very impressive (but easy to cross) edge. You will reach the Tuca de Posets (Posets Peak) (3,375 m.) after 6 hours, 4 and a half hours from the refuge.

High Mountain Excursion. It is essential that you be accompanied by a High Mountain Guide (if you do not have previous experience), wear mountain boots and material such as crampons and ice axe. Not recommended for children.